OpenStudy (anonymous):

Consider the quadratic function y=x^2-8x+195. How do I write the function in vertex form? How do I identify the vertex?

OpenStudy (jhannybean):

Completing the square method would work here.

OpenStudy (anonymous):

I have to complete the square also, but I still have to write it in vertex form and I don't know how to

OpenStudy (anonymous):

You start by completing the square

OpenStudy (anonymous):

how do I do that

OpenStudy (jhannybean):

Vertex form of the quadratic is: \(y=a(x-h)^2 +k\)

OpenStudy (jhannybean):

So we start by grouping, put a parenthesis around the x-terms

OpenStudy (perl):

h = -b/(2a) k = f(-b/(2a))

OpenStudy (perl):

and then plug in (0, 195) to get 'a'

OpenStudy (anonymous):

What is f?

OpenStudy (jhannybean):

\[y=(x^2-8x)+195\]\[c=\left(\frac{b}{2}\right)^2 = \left(\frac{-8}{2}\right)^2 = (-4)^2 = 16\]\[y=(x^2-8x+16)+195-16\]\[y=(x-4)^2+179\]

OpenStudy (jhannybean):

Basically, when you are completing the square, you are "completing" a quadratic inside the parenthesis. You find a `new` c value that is added to both the quadratic and the constant outside the quadratic, respectively.

OpenStudy (anonymous):

how do I find out what the vertex is?

OpenStudy (jhannybean):

\[y=a(x-\color{blue}h)^2 +\color{blue}k\]\[y=1(x-\color{blue}4)^2+\color{blue}{179}\]\[\text{vertex of parabola} = (h,k)\]

OpenStudy (jhannybean):

Make sense?

OpenStudy (anonymous):

so it would be (4, 179)? Not -4?

OpenStudy (anonymous):

how would I know if it has any x intercepts?

OpenStudy (jhannybean):

Yes, (4,179)

OpenStudy (anonymous):

Ok, I do understand how to do that now...how would I know if it has any x intercepts?

OpenStudy (jhannybean):

\[(x-4)^2+179=0\] Solve for x

OpenStudy (anonymous):

Thank you very much

OpenStudy (jhannybean):

No problem :)